# A new source of purely finite matricial fields

### 30 Mar 2026 v3

@@ Abstract.
A countable group $G$ is said to be matricial field (MF) if it admits a strongly converging sequence of approximate homomorphisms into matrices; i.e, the norms of polynomials converge to those in the left regular representation. $G$ is then said to be purely MF (PMF) if this sequence of maps into matrices can be chosen as actual homomorphisms. $G$ is further said to be purely finite field (PFF) if the image of each homomorphism is finite. By developing a new operator algebraic approach to these problems, we are able to prove the following result bringing several new examples into the fold. Suppose $G$ is a MF (resp., PMF, PFF) group and $H<G$ is separable (i.e., $H=\cap_{i\in\mathbb{N}}H_{i}$ where $H_{i}<G$ are finite index subgroups) and $K$ is a residually finite MF (resp., PMF, PFF) group. If either $G$ or $K$ is exact, then the amalgamated free product $G*_{H}(H\times K)$ is MF (resp., PMF, PFF).
Our work has several applications. Firstly, as a consequence of MF, the Brown–Douglas–Fillmore semigroups of many new reduced $C^{*}$-algebras are not groups. Secondly, we obtain that arbitrary graph products of residually finite exact MF (resp., PMF, PFF) groups are MF (resp., PMF, PFF), yielding a significant generalization of the breakthrough work of Magee–Thomas. Thirdly, our work resolves the open problem of proving PFF for fundamental groups of closed hyperbolic 3-manifolds. More generally all groups that virtually embed into RAAGs are PFF. Prior to our work, PFF was not known even in the case of free products. Our results are of geometric significance since PFF is the property that is used in Antoine Song’s approach towards the existence of minimal surfaces of negative curvature in round spheres.
@@ 1.1. Matricial field
Outside of intrinsic interest in random matrix theory, these properties have phenomenal applications. MF has applications to operator algebras. If a $C^{*}$-algebra is MF but not quasidiagonal, then $A$ has an extension by the compact operators $\mathbb{K}$ which is not invertible in the sense of Brown, Douglas and Fillmore [9] (see for instance the remark at the end of Section 2 in [44]). In particular if a group is non amenable (hence not quasidiagonal [22]) and MF, then $Ext(C^{*}_{r}(G))$ is not a group. PMF has certain geometric applications [36], including spectral gaps of certain Laplacians acting on vector bundles. PFF has remarkably found use in the theory of minimal surfaces through work of A. Song in his approach [48]. PFF is also interestingly in the spirit of soficity [39] for algebras: it is not just a finite-dimensional approximation of the reduced group $C^{*}$-algebra, but even a finite approximation. PPF is the most desirable property here not only because it encompasses all of the above, but also because it plays a fundamental role in deep applications to the study of optimal spectral gaps of graphs and hyperbolic manifolds [7, 28]. PPF is however quite difficult to access, especially because it fails in certain simple examples (see Subsection 1). One of the main novelties of the present article is that, despite this, we demonstrate that it is still possible in significant generality to bypass PPF and access PFF instead. For further details and references concerning these notions we point the reader to the beautiful surveys [37, 50].
@@
Importantly, these properties are much harder to prove than weaker representability properties such as soficity or Connes embeddability [39], because it involves establishing strong convergence. Despite the vast collection of examples of soficity and hyperlinearity, at the moment only a handful of examples of groups are known to be MF/PMF/PFF/PPF. Moreover, it is noteworthy to point out that proving such properties in each of these cases has been profoundly difficult and needed deep insights. The current list of examples includes MF for amenable groups [49]; PPF for free groups [7]; PPF for limit groups [34]; PMF for right angled Artin groups (RAAGs) [36]; MF for crossed products by free groups acting on amenable groups [40]; MF for $G_{1}*_{H}G_{2}$ where $G_{1},G_{2}$ are amenable and $H$ is a normal subgroup [43]. These properties are easily seen to be closed under taking subgroups and MF/PMF/PFF are closed under taking finite index overgroups (Lemma 7.1 in [34], see also Lemma 2.4). Being closed under finite index overgroups seems open in the case of PPF. It is currently known that MF/PMF is also closed under free products [12, 25]. Moreover, MF/PMF/PFF is also closed under direct products provided one of the groups is exact [21]. Strikingly, PPF is necessarily not closed under direct products, see [37] for a proof that $\mathbb{F}_{2}\times\mathbb{F}_{2}\times\mathbb{F}_{2}$ is not PPF. Interestingly, it is also known that $SL_{4}(\mathbb{Z})$ is not PMF [35]. Despite all of these results, the situation for wider natural families of groups, in particular for amalgamated free products, has remained a challenging open problem.
@@ 1.2. Main results
In this paper we are able to make significant new progress and expand on the collection MF, PMF and notably PFF groups by developing a new $C^{*}$-algebraic approach to the problem. Our main result is the following:
@@
We point out that strikingly even the case of free products remained open for the PFF property until our work. Outside of families of LERF groups (see a list of such examples in [19]), a nice instance of such inclusions $H<G$ which can be fed into our Corollary 1.2 is quasi-convex subgroups of cubulated hyperbolic groups. Indeed, cubulated hyperbolic groups are subgroups of RAAGs [1] and hence PFF (our Corollary 1.3 below), and quasi-convex subgroups are separable [23]. Another natural setting for separability is in the world of graph products [30]. The following is our next main Corollary:
@@
Our result above offers a significant generalization of the breakthrough work of Magee–Thomas [36]. We emphasize that this combination result yields many new examples of MF and PMF groups, and even in the case of RAAGs, our proof of PMF is new and succinct. More importantly, the reach of our results is higher and can handle the PFF property for RAAGs and therefore all groups that embed virtually in them. For fundamental groups of closed hyperbolic 3-manifolds, this has been an open problem in the field [^1] We thank Ramon van Handel for sharing with us the statement of this open problem., which is now settled. This particular result of ours is of geometric significance because PFF (for free and surface groups) is exactly the ingredient that goes into the breakthrough work of A. Song on minimal surfaces in Euclidean unit spheres [48]. A. Song’s paper accesses PFF via the PPF result of Bordenave–Collins [7], and the follow up paper [3] accesses PFF via Louder–Magee’s PPF for surface groups [34]. However, our results open the door to large new families of groups without having to go through PPF, which has obstructions as we shall describe soon. This naturally suggests generalizations of the works [48, 3]. We document the following below (see [1, 23] and also [4, Theorem 5.27]):
@@ 1.3. Comments to the reader
We would like to first discuss optimality aspects of our work. First note that exactness is a rather essential ingredient, in for instance Corollary 1.3. Indeed even in the tensor product case it is not known if MF is preserved in full generality [21]. A natural question arises of whether our results can extend to accessing PPF. We point out that this is in fact not possible. For instance, our Corollary 1.3 simply cannot generalize to PPF as stated because as mentioned earlier, Magee has shown that the group $\mathbb{F}_{2}\times\mathbb{F}_{2}\times\mathbb{F}_{2}$ is not PPF [37], and this is indeed a RAAG. For this reason PFF seems to be the best one can hope for in these general families of groups. The challenge with PPF is that it is not amenable to being stable under direct products (or finite index overgroups), due to the following fundamental issue: tensoring does not preserve the image being in the orthogonal complement of the invariant vectors for permutation matrices. Another point we would like to make is that it is at the moment an intractable open problem whether soficity or Connes embeddability is closed under amalgamated free products. In fact, considering the examples in Theorem 1.1, it is unclear whether the separability assumption can be dropped, simply because even Connes embeddability is not known in that generality.
@@
Insights into the proof
A key accomplishment of our work is to isolate a precise and potent connection between some of the deep tools going into the emerging program on selflessness in $C^{*}$-algebras, and strong convergence of unitary representations. In effect, our approach is able to entirely reduce the problem of constructing strongly converging matrix models of groups of the form $G*_{H}(H\times\mathbb{Z})$, to just the case of free groups $\mathbb{F}_{n}$ ([20, 7]) and elementary constructions such as induced representations and tensor models via exactness. On one hand, handling such amalgamated free product groups allows us to cover large families at the same time thanks to beautiful results in geometric group theory around separability, RAAGs and virtual specialness [1, 23, 57, 30]. On the other hand, this construction is precisely what allows us access to the deep $C^{*}$-correspondence theory machinery (see Theorem 2.2), in order to upgrade our maps from just the groups to the reduced $C^{*}$-algebras.
Firstly, our maps from Lemma 2.3 are built into an ultraproduct of amalgamated free products (by passing to normal cores from the separability assumptions) which are each by design virtually free, and therefore admit various strongly converging models. In order for our maps to preserve norms, we use a Fell’s absorption trick by tensoring with a copy of $G$ throughout the procedure. Interestingly, this is the reason why we need exactness in the case that we replace $\mathbb{Z}$ with an arbitrary residually finite group $L$, and also the reason why one cannot automatically use these arguments to build PPF models. Continuing with the argument, we crucially use a remarkable isomorphism theorem (Theorem 2.2), which says that a universal Toeplitz–Pimnser type algebra satisfying just some simple orthogonality conditions, is automatically isomorphic to a reduced amalgamated free product of the type we have been considering. All we need to do is check the prescribed equalities with the “creation operators”, and we are then in business for strong convergence. We strongly urge the reader who is not familiar with these $C^{*}$-algebraic considerations to consider reading our proof in parallel with Section 4.6 of [10], which has a very neat treatment of the breakthroughs in $C^{*}$-correspondence theory around Fowler–Muhly–Raeburn’s gauge invariant-uniqueness theorem [16] and its consequences including the isomorphism theorem we use here.
@@ 1.4. Acknowledgments
We thank M. Junge for an insightful question that stimulated this work during the workshop “Graduate mini-school on sofic groups and Connes embedding problem” at UT Austin during March 18–22, 2026. We thank UT Austin and the organizers for providing a stimulating environment to work in. We are indebted to R. van Handel for several long correspondences and his extremely generous feedback and support. We thank D. Jekel and N. Ozawa for helpful comments and corrections. We also thank B. Hayes and F. Fournier-Facio for helpful conversations regarding references and M. Linton for his suggestion of a quick proof of Lemma 2.5. The second author is grateful to U. Bader for his encouragement and helpful conversations. This work was supported partially by the second author’s NSF grant DMS 2350049. The fourth author was supported by the Engineering and Physical Sciences Research Council (UK), grant EP/X026647/1.
@@ 2. Proof of main result
We first prove the case $L=\mathbb{Z}$. As in Lemma 2.3, we set $K_{i}=\bigcap_{g\in G}gH_{i}g^{-1}$, the normal core of $H_{i}$. Note that $K_{i}$ is of finite index in $G$, so we have $G/K_{i}\ast_{H_{i}/K_{i}}(H_{i}/K_{i}\times\mathbb{Z})$ is PFF by Lemma 2.5. These are exact since amalgamated free products of exact groups are exact [13, Theorem 3.2]. Since $G$ is MF (resp., PMF, PFF) by hypothesis, we have that
@@
is also MF (resp., PMF, PFF) by [21, Prop 3.2] or Lemma 2.4. Now the result follows from Lemma 2.3.

### 02 Apr 2026 v4

@@ Abstract.
A countable group $G$ is said to be matricial field (MF) if it admits a strongly converging sequence of approximate homomorphisms into matrices; i.e, the norms of polynomials converge to those in the left regular representation. $G$ is purely MF (PMF) if these maps are actual homomorphisms, and $G$ is further purely finite field (PFF) if the image of each homomorphism is finite. By developing a new operator algebraic approach to these problems, we are able to prove the following result bringing several new examples into the fold. Suppose $G$ is a MF (resp., PMF, PFF) group and $H<G$ is separable (i.e., $H=\cap_{i\in\mathbb{N}}H_{i}$ where $H_{i}<G$ are finite index subgroups) and $K$ is a residually finite MF (resp., PMF, PFF) group. If either $G$ or $K$ is exact, then the amalgamated free product $G*_{H}(H\times K)$ is MF (resp., PMF, PFF). Our work has several applications, we list some below:
  1. (1)
    The Brown–Douglas–Fillmore semigroups of many new examples of reduced group $C^{*}$-algebras are shown to be not groups.
  2. (2)
    Arbitrary group doubles $G*_{H}G$ of MF (resp., PMF, PFF) over separable subgroups $H$ are MF (resp., PMF, PFF). Moreover, GHG*H is PFF whenever $G,H$ are PFF, and either $G$ or $H$ is exact.
  3. (3)
    Arbitrary graph products of residually finite exact MF (resp., PMF, PFF) groups are MF (resp., PMF, PFF), yielding a significant generalization of the breakthrough work of M. Magee and J. Thomas.
  4. (4)
    The open problem of proving PFF for fundamental groups of closed hyperbolic 3-manifolds is resolved. This has geometric significance in the theory of minimal surfaces via A. Song’s approach.
@@ 1.1. Matricial field
Outside of intrinsic interest in random matrix theory, these properties have phenomenal applications. MF has applications to operator algebras. If a $C^{*}$-algebra is MF but not quasidiagonal, then $A$ has an extension by the compact operators $\mathbb{K}$ which is not invertible in the sense of Brown, Douglas and Fillmore [9] (see for instance the remark at the end of Section 2 in [49]). In particular if a group is non amenable (hence not quasidiagonal [27]) and MF, then $Ext(C^{*}_{r}(G))$ is not a group. PMF has certain geometric applications [41], including spectral gaps of certain Laplacians acting on vector bundles. PFF has remarkably found use in the theory of minimal surfaces through work of A. Song in his approach [53]. PFF is also interestingly in the spirit of soficity [44] for algebras: it is not just a finite-dimensional approximation of the reduced group $C^{*}$-algebra, but even a finite approximation. PPF is the most desirable property here not only because it encompasses all of the above, but also because it plays a fundamental role in deep applications to the study of optimal spectral gaps of graphs and hyperbolic manifolds [7, 33]. PPF is however quite difficult to access, especially because it fails in certain simple examples. One of the main novelties of the present article is that we demonstrate that it is still possible in significant generality to bypass PPF and access PFF instead. For further details and references concerning these notions we point the reader to the beautiful surveys [42, 56].
@@
Importantly, these properties are much harder to prove than weaker representability properties such as soficity or Connes embeddability [44], because it involves establishing strong convergence. Despite the vast collection of examples of soficity and hyperlinearity, at the moment only a handful of examples of groups are known to be MF/PMF/PFF/PPF. Moreover, it is noteworthy to point out that proving such properties in each of these cases has been profoundly difficult and needed deep insights. The current list of examples includes MF for amenable groups [54]; PPF for free groups [7] (see also [25, 12]); PPF for limit groups [39]; PMF for right angled Artin groups (RAAGs) [41] (see also [13]); MF for crossed products by free groups acting on amenable groups [45]; MF for $G_{1}*_{H}G_{2}$ where $G_{1},G_{2}$ are amenable and $H$ is a normal subgroup [48]. These properties are easily seen to be closed under taking subgroups and MF/PMF/PFF are closed under taking finite index overgroups (Lemma 7.1 in [39], see also Lemma 2.4). Being closed under finite index overgroups seems open in the case of PPF. It is currently known that MF/PMF is also closed under free products [15, 30]. Moreover, MF/PMF/PFF is also closed under direct products provided one of the groups is exact [26]. Strikingly, PPF is necessarily not closed under direct products, see [42] for a proof that $\mathbb{F}_{2}\times\mathbb{F}_{2}\times\mathbb{F}_{2}$ is not PPF (an identical argument of Magee shows $G\times H\times K$ is not PPF when $G,H$ are nonamenable and $K$ is nontrivial). Interestingly, it is also known that $SL_{4}(\mathbb{Z})$ is not PMF [40]. Despite all of these results, the situation for wider natural families of groups, in particular for amalgamated free products, has remained a challenging open problem.
@@ 1.2. Main results
In this paper we are able to make significant conceptual progress and expand on the collection MF, PMF and notably PFF groups by developing a new $C^{*}$-algebraic approach to the problem. Our main result is the following:
@@
We point out that strikingly the case of free products had remained open for the PFF property until our result above. It turns out that conjugation by random unitaries as in [15] destroys the finiteness of the representations which is crucial in PFF. Note also that the exactness assumption on one of the two groups is still necessary to apply our Theorem 1.1 in this situation, via passing to direct products. In the context of the separability assumptions in Theorem 1.1 and Corollary 1.2, we remark that outside of the families of LERF groups (see a list of such examples in [24]), a nice instance of separable inclusions $H<G$ which can be fed into our results is quasi-convex subgroups of cubulated hyperbolic groups. Indeed, cubulated hyperbolic groups are subgroups of RAAGs [1] and hence PFF (our Corollary 1.3 below), and quasi-convex subgroups are separable [28]. Another natural setting for separability is in the world of graph products [35]. The following is our next main Corollary:
@@
Our result above offers a significant generalization of the breakthrough work of Magee–Thomas [41]. We emphasize that this combination result yields many new examples of MF and PMF groups, and even in the case of RAAGs, our proof of PMF is new and succinct. More importantly, the reach of our results is higher and can handle the PFF property for RAAGs and therefore all groups that embed virtually in them. For fundamental groups of closed hyperbolic 3-manifolds, this has been an open problem in the field [^1] We thank Ramon van Handel for sharing with us the statement of this open problem., which is now settled. This particular result of ours is of geometric significance, as we shall explain soon, because PFF (for free and surface groups) is exactly the ingredient that goes into the breakthrough work of A. Song on minimal surfaces in Euclidean unit spheres [53]. A. Song’s paper accesses PFF via the PPF result of Bordenave–Collins [7], and the follow up paper [3] accesses PFF via Louder–Magee’s PPF for surface groups [39]. However, our results open the door to large new families of groups without having to go through PPF. We document our result below (see [1, 28] and also [4, Theorem 5.27]):
@@
We are grateful to A. Song for sharing his helpful insights regarding the potential of the above result for geometry, which we briefly explain for the benefit of the reader before concluding this section. Given a closed manifold $M$ with fundamental group $G$, and given an $n$-dimensional unitary representation $\rho_{n}$ of $G$ such that $\rho_{n}(G)$ is finite, there is a natural way to construct a corresponding closed minimal submanifold $\Sigma_{n}$ inside some unit sphere. Finiteness, granted by the PFF property, is crucial here since the objects of interest are closed submanifolds. When such a construction is fully justified, it turns out that PFF implies convergence of the shape of $\Sigma_{n}$ to a limit shape which can be deciphered, thanks to a good understanding of the regular representation. In the case of dimension $2$, closed 2-dimensional minimal surfaces that are almost hyperbolic have been constructed very recently in this manner [53, 3], answering old problems of S. T. Yau. Our Corollary 1.4 suggests strongly that this phenomena should carry over in dimension $3$ for instance, and more generally whenever $G$ virtually embeds in a RAAG. It is plausible that there are many closed minimal submanifolds $\Sigma^{d}$ of dimension $d\geq 3$ inside Euclidean spheres, whose limit geometry is that of a hyperbolic $d$-space. This would be satisfying from a geometric viewpoint as it would give the existence, in any dimension, of large families of closed minimal submanifolds in spheres whose limit shape is completely understood and canonical.
@@ 1.3. Comments to the reader
We would like to first discuss optimality aspects of our work. First note that exactness is a rather essential ingredient in certain cases, for instance Corollary 1.3. Indeed even in the tensor product case it is not known if MF is preserved in full generality [26]. A natural question arises of whether our main result and its corollaries can extend to accessing PPF. We point out that this is in fact not possible. For instance, our Corollary 1.3 simply cannot generalize to PPF as stated because as mentioned earlier, Magee has shown that the group $\mathbb{F}_{2}\times\mathbb{F}_{2}\times\mathbb{F}_{2}$ is not PPF [42], and this is indeed a RAAG. In fact, Magee’s argument shows that $G\times H\times K$ is not PPF whenever $G,H$ are nonamenable and $K$ is nontrivial. The real challenge with PPF is that it is not amenable to being stable under direct products (or finite index overgroups), due to the following fundamental issue: tensoring does not preserve the image being in the orthogonal complement of the invariant vectors for permutation matrices.
Another remark we would like to make is that it is at the moment an intractable open problem whether soficity or Connes embeddability is closed under amalgamated free products. In fact, considering the examples in Theorem 1.1, it is unclear whether the separability assumption can be dropped, simply because even Connes embeddability is not known in that generality.
@@
Insights into Theorem 1.1
A key accomplishment of our work is to isolate a precise and powerful connection between some of the deep tools going into the emerging program on selflessness in $C^{*}$-algebras, and strong convergence of unitary representations. In effect, our approach is able to entirely reduce the problem of constructing strongly converging matrix models of groups of the form $G*_{H}(H\times\mathbb{Z})$, to just the case of free groups $\mathbb{F}_{n}$ ([25, 7]) and elementary constructions such as induced representations and tensor models via exactness. On one hand, handling such amalgamated free product groups allows us to cover large families at the same time thanks to beautiful results in geometric group theory around separability, RAAGs and virtual specialness [1, 28, 63, 35]. On the other hand, this construction is precisely what allows us access to the deep $C^{*}$-correspondence theory machinery (see Theorem 2.2), in order to upgrade our maps from just the groups to the reduced $C^{*}$-algebras.
Firstly, our maps from Lemma 2.3 are built into an ultraproduct of amalgamated free products (by passing to normal cores from the separability assumptions) which are each by design virtually free, and therefore admit various strongly converging models. In order for our maps to preserve norms, we use a Fell’s absorption trick by tensoring with a copy of $G$ throughout the procedure. Interestingly, this is the reason why we need exactness in the case that we replace $\mathbb{Z}$ with an arbitrary residually finite group $L$, and also the reason why one cannot automatically use these arguments to build PPF models. Continuing with the argument, we crucially use a remarkable isomorphism theorem (Theorem 2.2), which says that a universal Toeplitz–Pimnser type algebra satisfying just some simple orthogonality conditions, is automatically isomorphic to a reduced amalgamated free product of the type we have been considering. All we need to do is check the prescribed equalities with the “creation operators”, and we are then in business for strong convergence.
We would like to remark that the usage of isomorphism theorems of the above nature stand on tradition in operator algebras, as explained in Section 4.6 of [10]. For instance, it is used crucially in the proof of permanence of exactness for reduced amalgamated free products [17], and additionally weak exactness [55]. A universality statement in the context of Toeplitz algebras associated to RAAGs, due to Crisp and Laca [16], is also employed (albeit in combination with random matrix theory methods [14]) in the PMF work [42]. Our current work reveals in hindsight that this was in fact a potent early insight of Magee–Thomas.
Note to the reader
We strongly suggest the reader who is not familiar with the $C^{*}$-algebraic considerations herein to read our proof in parallel with Section 4.6 of [10], which has a very neat treatment of the breakthroughs in $C^{*}$-correspondence theory around Fowler–Muhly–Raeburn’s gauge invariant-uniqueness theorem [21] and its consequences including the isomorphism theorem we use here.
@@ 1.4. Acknowledgments
We thank M. Junge for an insightful question that stimulated this work during the workshop “Graduate mini-school on sofic groups and Connes embedding problem” at UT Austin during March 18–22, 2026. We thank UT Austin and the organizers for providing a stimulating environment to work in. We are indebted to R. van Handel for several long correspondences and his extremely generous feedback and support. We are grateful to A. Song for his encouragement, and generous feedback. We thank D. Jekel and N. Ozawa for helpful comments and corrections. We also thank B. Hayes and F. Fournier-Facio for helpful conversations regarding references and M. Linton for his suggestion of a quick proof of Lemma 2.5. The second author is grateful to U. Bader for his encouragement and helpful conversations. This work was supported partially by the second author’s NSF grant DMS 2350049. The fourth author was supported by the Engineering and Physical Sciences Research Council (UK), grant EP/X026647/1.
@@ 2. Proof of main result
We first prove the case $L=\mathbb{Z}$. As in Lemma 2.3, we set $K_{i}=\bigcap_{g\in G}gH_{i}g^{-1}$, the normal core of $H_{i}$. Note that $K_{i}$ is of finite index in $G$, so we have $G/K_{i}\ast_{H_{i}/K_{i}}(H_{i}/K_{i}\times\mathbb{Z})$ is PFF by Lemma 2.5. These are exact since amalgamated free products of exact groups are exact [18, Theorem 3.2]. Since $G$ is MF (resp., PMF, PFF) by hypothesis, we have that
@@
is also MF (resp., PMF, PFF) by [26, Prop 3.2] or Lemma 2.4. Now the result follows from Lemma 2.3.